OpenStudy (anonymous):

What is the solution of the equation? Eliminate any extraneous solutions (-2x+6)1/5 = (-8+10x)1/5

OpenStudy (austinl):

\((-2x+6)\dfrac{1}{5} = (-8+10x)\dfrac{1}{5}\) First we can cancel out the \(\dfrac{1}{5}\)'s \((-2x+6) = (-8+10x)\) Get variables on one side and constants on the other.

OpenStudy (anonymous):

okay once i do that, what happens next

OpenStudy (austinl):

What would that look like? Then after you get that done, you would divide the constants by the coefficient of the variable.

OpenStudy (anonymous):

ahh okay gotcha, i got it from here, thanks man

OpenStudy (austinl):

No problemo!

OpenStudy (anonymous):

what was the correct answer?